OpenStudy (anonymous):

Andrea has 1 hour to spend training for an upcoming race. She completes her training by running full speed the distance of the race and walking back the same distance to cool down. If she runs at a speed of 7mph and walks back at a speed of 3 mph, how long should she plan to spend walking back?

OpenStudy (anonymous):

is this all the info?

OpenStudy (amistre64):

d = s*t its a ratio of 7/3, or is it 3/7?

OpenStudy (anonymous):

yeah this is a;ll the information

OpenStudy (amistre64):

d = 7n d = 3m 7n = 3m m = 7n/3 ; so it seems to depend on another bit of info if i see it right

OpenStudy (amistre64):

maybe we derive the distance by equating the full time out to be an hour and divvy the time between run and walk?

OpenStudy (anonymous):

this is a brain tingler is there anyway we can use ratios?

OpenStudy (anonymous):

i dont think so.... it is just asking for the number of hours

OpenStudy (amistre64):

assume we know the distance; say 528 miles; how would it solve then?

OpenStudy (amistre64):

... 28 miles lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

mmm we cooould

OpenStudy (amistre64):

get a particular solution and work it into a general :)

OpenStudy (amistre64):

7*3 = 21; so try that :)

OpenStudy (anonymous):

okay wait i got it

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

the speed ratio is 7:3

OpenStudy (anonymous):

so shouldnt the time also be 7:3

OpenStudy (anonymous):

or 3:7?

OpenStudy (anonymous):

60 = m + n 7 * m = 3 * n solve this

OpenStudy (amistre64):

that looks plausible

OpenStudy (anonymous):

i have no idea... i am lost

OpenStudy (amistre64):

18 min running and 42 minutes walking sounds good :)

OpenStudy (anonymous):

60-n = m 7 * m = 3n 7(60-n) = 3 n 420-7n = 3n 420 = 3n+7n 420 = 10n 42 = n m = 60-42 = 18 7 * 18/60 = 2.1 miles 3 * 42/60 = 2.1 miles

OpenStudy (anonymous):

there should be it

OpenStudy (anonymous):

yeah thats what i was thinking. u agree amistre?

OpenStudy (amistre64):

i agree :)

OpenStudy (anonymous):

sweets. how is that kelly?

OpenStudy (anonymous):

so the correct answer is 2.1 miles?

OpenStudy (amistre64):

lol ... distance isnt time

OpenStudy (anonymous):

arent you looking for time?

OpenStudy (anonymous):

its 42 minutes walking back

OpenStudy (amistre64):

time running = 18 and time walking = 42 distance of race = 2.1 maybe

OpenStudy (anonymous):

whoops.... i got confused... it is asking for hours

OpenStudy (anonymous):

well, it the question it states that she have one hour to practise in, so the answer should be 42 minutes as far as i can see.